# ARES: Alternating Reinforcement Learning and Supervised Fine-Tuning for Enhanced Multi-Modal Chain-of-Thought Reasoning Through Diverse AI Feedback

### v1

@@ Abstract
Large Multimodal Models (LMMs) excel at comprehending human instructions and demonstrate remarkable results across a broad spectrum of tasks. Reinforcement Learning from Human Feedback (RLHF) and AI Feedback (RLAIF) further refine LLMs by aligning them with specific preferences. These methods primarily use ranking-based feedback for entire generations. With advanced AI models (Teacher), such as GPT-4 and Claude 3 Opus, we can request various types of detailed feedback that are expensive for humans to provide. We propose a two-stage algorithm ARES that Alternates REinforcement Learning (RL) and Supervised Fine-Tuning (SFT). First, we request the Teacher to score how much each sentence contributes to solving the problem in a Chain-of-Thought (CoT). This sentence-level feedback allows us to consider individual valuable segments, providing more granular rewards for the RL procedure. Second, we ask the Teacher to correct the wrong reasoning after the RL stage. The RL procedure requires massive efforts for hyperparameter tuning and often generates errors like repetitive words and incomplete sentences. With the correction feedback, we stabilize the RL fine-tuned model through SFT. We conduct experiments on multi-model dataset ScienceQA and A-OKVQA to demonstrate the effectiveness of our proposal. ARES rationale reasoning achieves around 70% win rate against baseline models judged by GPT-4o. Additionally, we observe that the improved rationale reasoning leads to a 2.5% increase in inference answer accuracy on average for the multi-modal datasets. 111Code: https://github.com/Amyyyyeah/ARES
@@ 1 Introduction
Large Language Models (LLMs) and Large Multimodal Models demonstrate remarkable performance across diverse language and multi-modal tasks (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Zhang et al., 2022a; Liu et al., 2023a; Goyal et al., 2023). However, these Large Models (LMs) often generate toxic and biased content (Gehman et al., 2020; Tamkin et al., 2021) because LMs are primarily trained to predict the next token based on extensive corpus datasets. To align LM behavior more closely with user preferences, previous works (Glaese et al., 2022; Ouyang et al., 2022) fine-tune their models using Reinforcement Learning from Human Feedback (RLHF). Furthermore, with advancements in LMs, an advanced LM feedback start replacing costly human feedback, yielding Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023; Bai et al., 2022; Yuan et al., 2024).
However, RLHF and RLAIF encounter two significant challenges. First, both methods often utilize ranking-based feedback (Ouyang et al., 2022), which orders the preferences of generated samples. For instance, if sample is preferred over sample , the model is fine-tuned to generate more outputs like and fewer like . However, if contains certain valuable parts, these parts are often disregarded. To alleviate this issue, Lightman et al. (2023); Luo et al. (2024) propose sentence-level feedback, applying it solely to the reward model without Reinforcement Learning (RL), called the Process-supervised Reward Model (PRM). It demonstrates its potential through search algorithms on the PRM, such as best-of- or Monte Carlo Tree Search (MCTS). Furthermore, Wang et al. (2024) demonstrate the effectiveness of RL with sentence-level feedback by heuristically scoring each sentence in math problems, where evaluating the predicted answer is straightforward. Thus, sentence-level feedback exhibits significant promise compared to existing ranking-based feedback. Nonetheless, acquiring sentence-level feedback is more costly than ranking-based feedback.
@@
In this work, we aim to address the two challenges mentioned above through various types of feedback using an advanced AI model as a Teacher. Many advanced AI models, including GPT-4 and Claude 3 Opus, are already used as evaluators for many tasks and generate reliable human-level answers (Liu et al., 2023b; Sottana et al., 2023). We request a score from Teacher for each sentence, ranging from to . Each score indicates how much a sentence contributes to solving the problem. This provides detailed reward feedback to the training model and can be applied to both mathematical and more general multi-modal Chain-of-Thought (CoT) (or rationale reasoning) problems. We ask the Teacher to identify and correct minor errors in the RL results, such as incorrect or cut-off parts. With this corrected dataset, we fine-tune the model using Supervised Fine-Tuning (SFT). This stage allows the model to maximize the rewards while properly deviating from the pretraining distribution. In summary, We propose a hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning.
@@
ARES stabilizes an RL fine-tuned model through SFT and properly reflects the direction RL wants to change without massive hyperparameter tuning.
@@ 2 Methodology
This section briefly introduces the preliminaries in Section 2.1 and present our two-stage hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning. We request a score for each sentence in the Chain-of-Thought (CoT) from the advanced AI model (Teacher) to determine how much it contributes to solving the problem (Section 2.2). We perform Reinforcement Learning (RL) with the score feedback on our training model. Teacher corrects minor errors such as truncated or slightly incorrect sentences, thereby performing Supervised Fine-Tuning (SFT) (Section 2.2).
@@ 2.1 Preliminaries
For an input , a transformer based (Vaswani et al., 2023) model parameterized by generates the output composed of sentences .
where indicates previous tokens. To proceed with RL finetuning, Ouyang et al. (2022) train an outcome-supervised reward model (ORM) using ranking-based feedback. With more fine-grained feedback like sentence-level, Lightman et al. (2023); Wang et al. (2024) train a process-supervised reward model (PRM). Instead of training a reward model, we request score feedback for each sentence from an advanced AI such as GPT-4 where is previous sentences.
@@ 2.2 Reinforcement Learning
Reinforcement Learning (RL) fine-tunes our model to maximize sum of sentence rewards from an advanced AI such as GPT-4 and Claude 3 Opus. The RL objective is as follows:
@@
Sentence-Level Nuanced Feedback: We request a score between and for each sentence in CoT through the advanced AI for RL. The closer the score is to , the more relevant and helpful it is to solving the problem. Table 5 presents the prompt format. We additionally shift the reward distribution by to center it at (Zheng et al., 2023). Therefore, the actual range is from to . Using these nuanced scores, the RL fine-tuned model exhibits emergent behaviors (please refer to Section 4). This allows us to understand the direction in which the model wants to change through RL.
Advantages of Using Advanced AI for Score Feedback: Although calling the API has disadvantages, such as incurring costs or facing usage limits, there exist several advantages of using the advanced AI for feedback. First, there is no need to train a reward model. Second, as the RL fine-tuned model begins to generate out-of-distribution outputs that differ from the data used to train the reward model, it becomes challenging for the trained reward model to provide accurate rewards. However, this out-of-distribution problem is effectively addressed with the advanced AI.
@@ 2.3 Correction: Supervised Fine-Tuning
The RL fine-tuning procedure makes model changes to maximize the reward sum, such as correcting mistakes or explaining why other options cannot be the answer. However, without highly tuned hyperparameters (Eimer et al., 2023), the model after the RL phase may often result in errors such as repeated sentences, truncated sentences, or incorrect content for some data points. (See examples in Appendix D.)
Correction Feedback: Given the success of LLMs and LMMs in a wide range of areas (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022a), we do not need to be restricted to requesting feedback in the form of scores. We request correction feedback from advanced AI (Teacher) for sentences containing errors after the RL process, and obtain a corrected dataset . Since the supervised fine-tuning is more stable and finding appropriate hyperparameters is easier than RL, we proceed with supervised fine-tuning using exactly as in common autoregressive model (Vaswani et al., 2023) training to stabilize the RL fine-tuned model. This reduces the burden of RL’s exhaustive hyperparameter tuning and properly guides the direction in which the training model wants to change.
How Correction Feedback Helps RL: RL basically increases the probability of positively rewarded actions (or sentences) and decreases the probability for negative rewards. The direction of learning is determined by the reward (scalar) value. However, the opposite direction of the reward is sometimes required. For example, suppose there is a truncated sentence in CoT. gets a negative score because it is an incomplete sentence (Table 11). If there is no correction stage, the probability of is simply reduced. What if contains some valuable part? This valuable part is ignored, and its probability decreases. To alleviate this issue, we instead receive the corrected sentence as feedback and encourage the training model to generate complete sentences, which is very challenging to achieve with only RL.
@@
2.4 Algorithm Detail
We propose a hybrid algorithm Alternating between REinforcement learning and Supervised fine-tuning (ARES). Figure 2 illustrates the ARES pipeline. First, we prepare a model with a given training dataset and generate rationale reasoning composed of several sentences for input. For the RL procedure to align the training model with a preference, we request scores for each sentence. The RL result may include some incorrect parts (colored as the red 4th sentence in the RL result box), but it aims to maximize the rewards provided. Next, we request correction feedback and create a corrected dataset (colored as the blue 3rd sentence in the Corrected Rationale box) for the supervised fine-tuning stage. We then repeat the process from the RL stage until convergence.
@@ 3 Experimental Setup
Baselines: We mainly compare our method with Multimodal-CoT (MM-CoT) (Zhang et al., 2023b) as the baseline because it utilizes reasoning chains to solve multi-modal tasks. MM-CoT leverages two distinct models: the first generates rationale for a given problem, and the second, an inference model, takes the concatenated input (problem and generated rationale). This separated framework shows improved performance, even for relatively small models such as (Chia et al., 2023) (M) and (M). We use the rationale model provided by MM-CoT for ScienceQA and retrain the rationale model ourselves for A-OKVQA because there is no provided model.
Prompts for Feedback: Since our proposed ARES requests different types of feedback for each stage, a corresponding prompt exists separately. We use Claude 3 Haiku for all training to get feedback because it is about times cheaper than the top competing models, yet still demonstrates decent performance. We first request scores ranging from to for each sentence in CoT to proceed with the RL stage. To obtain reasonable scores, we let Haiku consider the starting point of thought, the process of elimination, or true statements. (See Table 5.)
@@
Training Details: For the RL stage, we use a learning rate of and epochs for PPO with a batch size of for both ScienceQA and A-OKVQA. The learning rate for is with epochs for PPO and a batch size of for both tasks. We proceed with 2 rounds of our pipeline for and 2 rounds for for ScienceQA. For A-OKVQA, we proceed with 1 round for both model sizes. For the SFT stage for correction, we follow the hyperparameters used in MM-CoT for both model sizes. Additionally, we replace MM-CoT’s inference model, which is the same size as the rationale model, with a LoRA adapter added to the rationale model (Figure 4). The LoRA adapter effectively refers to the rationale model’s features with a small number of weights and infers answers. Please refer to the more detailed settings in Appendix C.
@@ 4.2 Guide RL with Correction
Despite the benefits of RL, hyperparameter tuning often requires massive effort. Without meticulous tuning, the RL fine-tuned model may produce errors such as repetitive or incomplete sentences. To address these issues, we include a supervised fine-tuning (SFT) stage after RL to correct these errors. SFT is more stable than RL. We evaluate how well the SFT stage corrects errors caused by the RL stage for various RL hyperparameters. We test various RL hyperparameters such as learning rate = {5e-6, 1e-5, 2e-5, 5e-5}, batch size = {2, 4, 8, 16, 32}, and PPO epoch = {5, 10, 15}. As a result of RL, we observe that some of the sentences in rationale chains are repetitive or truncated (see Table 11 and 10). The SFT stage, with correction feedback, reflects what RL wants to achieve and appropriately guides it. However, excessive RL learning rates or epochs cause serious degeneration of the model, such as producing no output or generating strange words, and the results of correction feedback are also unreasonable.
@@ 4.4 Inference Accuracy
We investigate whether the improved rationale also contributes to answer inference accuracy. Table 2 shows the main results of answer inference on the ScienceQA. We evaluate our base model against the MM-CoT baseline. achieves a 2.79% improvement compared to the corresponding baseline (). The large model () shows some minimal improvement compared to the corresponding baseline. However, it’s worth noting that despite this seemingly small gain, beats to B LLaVA (Liu et al., 2023a). This minimal improvement may be due to the 9.5% of ScienceQA problems needing more rationale reasoning (around 9.5% problems have empty rationale reasoning). The RL stages can only eliminate some empty rationale reasoning, which requires numerous ARES pipeline rounds. Above all, our main goal is to assess how the RL stage works and how the SFT stage aids RL.
@@ 5 Related Work
Chain-of-Thought (CoT) is a multi-step reasoning method for problem-solving that encourages LLMs to consider the intermediate reasoning steps. Zero-Shot-CoT (Kojima et al., 2023) promotes CoT by using prompts such as "Let’s think step by step" for LLMs. For Few-Shot-CoT (Zhang et al., 2022b; Wei et al., 2023), a few examples with reasoning processes are provided, allowing the model to refer to these examples and understand how to perform CoT. Wei et al. (2023) reveal that this CoT technique positively impacts the performance of large models (B), but has minimal effect on smaller models. MM-CoT (Zhang et al., 2023b) suggest that CoT is beneficial even for relatively small models, such as 200M, if the model that generates intermediate reasoning and the model that infers the answer are separated. We find that simply adding a LoRA adapter (Hu et al., 2021) to the reasoning model shows comparable performance. This framework allows the LoRA adapter to effectively refer to all features from raw text to latent features to generate answers.

### v2

@@ Abstract
Large Multimodal Models (LMMs) excel at comprehending human instructions and demonstrate remarkable results across a broad spectrum of tasks. Reinforcement Learning from Human Feedback (RLHF) and AI Feedback (RLAIF) further refine LLMs by aligning them with specific preferences. These methods primarily use ranking-based feedback for entire generations. With advanced AI models (Teacher), such as GPT-4 and Claude 3 Opus, we can request various types of detailed feedback that are expensive for humans to provide. We propose a two-stage algorithm ARES that Alternates REinforcement Learning (RL) and Supervised Fine-Tuning (SFT). First, we ask the Teacher to score how much each sentence contributes to solving the problem in a Chain-of-Thought (CoT). This sentence-level feedback allows us to consider individual valuable segments, providing more granular rewards for the RL procedure. Second, we ask the Teacher to correct wrong reasoning after the RL stage. The RL procedure requires substantial hyperparameter tuning and often generates errors such as repetitive words and incomplete sentences. With correction feedback, we stabilize the RL fine-tuned model through SFT. We conduct experiments on the multi-modal datasets ScienceQA and A-OKVQA to demonstrate the effectiveness of our proposal. The ARES rationale achieves around 70% win rate compared to baseline models judged by GPT-4o. Additionally, we observe that the improved rationale reasoning leads to a 2.5% increase in inference answer accuracy on average for the multi-modal datasets. 111Code: https://github.com/Amyyyyeah/ARES
@@ 1 Introduction
Large Language Models (LLMs) and Large Multimodal Models demonstrate remarkable performance across diverse language and multi-modal tasks (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Zhang et al., 2022a; Liu et al., 2023a; Goyal et al., 2023). However, these Large Models (LMs) often generate toxic and biased content (Gehman et al., 2020; Tamkin et al., 2021) because LMs are primarily trained to predict the next token based on extensive corpus datasets. To align LM behavior more closely with user preferences, previous works (Glaese et al., 2022; Ouyang et al., 2022) fine-tune their models using Reinforcement Learning from Human Feedback (RLHF). Furthermore, with advancements in LMs, advanced LM feedback can replace costly human feedback, yielding Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023; Bai et al., 2022; Yuan et al., 2024).
However, RLHF and RLAIF encounter two significant challenges. First, both methods often utilize ranking-based feedback (Ouyang et al., 2022), which orders the generated samples by preference. For instance, if sample is preferred over sample , the model is fine-tuned to generate more outputs like and fewer like . However, if contains certain valuable parts, they may be discarded. To alleviate these issues, Lightman et al. (2023); Luo et al. (2024) propose sentence-level feedback, applying it solely to the reward model without Reinforcement Learning (RL), called the Process-supervised Reward Model (PRM). It demonstrates its potential through search algorithms on the PRM, such as best-of- or Monte Carlo Tree Search (MCTS). Furthermore, Wang et al. (2024) demonstrate the effectiveness of RL with sentence-level feedback by heuristically scoring each sentence in math problems, where evaluating the predicted answer is straightforward. Thus, sentence-level feedback exhibits significant promise compared to existing ranking-based feedback. Nonetheless, acquiring sentence-level feedback is more costly than ranking-based feedback.
@@
In this work, we aim to address the two challenges mentioned above through various types of feedback using an advanced AI model as a Teacher. Many advanced AI models, including GPT-4 and Claude 3 Opus, are already used as evaluators for many tasks and generate reliable human-level answers (Liu et al., 2023b; Sottana et al., 2023). We request a score from the Teacher for each sentence, ranging from to . Each score indicates how much a sentence contributes to solving the problem. This provides detailed reward feedback to the training model and can be applied to both mathematical and more general multi-modal Chain-of-Thought (CoT) (or rationale reasoning) problems. We ask the Teacher to identify and correct minor errors in the RL results, such as incorrect or cut-off parts. With this corrected dataset, we fine-tune the model using Supervised Fine-Tuning (SFT). This stage allows the model to maximize the rewards while properly deviating from the pretraining distribution. In summary, we propose a hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning.
@@
ARES properly reflects the direction in which the model is fine-tuned through RL and stabilizes the RL fine-tuning procedure with SFT.
@@ 2 Methodology
This section briefly introduces the preliminaries in Section 2.1 and presents our two-stage hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning. We request a score for each sentence in the Chain-of-Thought (CoT) from the advanced AI model (Teacher) to determine how much it contributes to solving the problem (Section 2.2). We perform Reinforcement Learning (RL) with the score feedback on our training model. The Teacher corrects minor errors such as truncated or slightly incorrect sentences, thereby performing Supervised Fine-Tuning (SFT) (Section 2.2).
@@ 2.1 Preliminaries
For an input , a Transformer-based (Vaswani et al., 2023) model parameterized by generates the output composed of sentences .
where indicates previous tokens. To proceed with RL finetuning, Ouyang et al. (2022) train an outcome-supervised reward model (ORM) using ranking-based feedback. With more fine-grained feedback like sentence-level, Lightman et al. (2023); Wang et al. (2024) train a process-supervised reward model (PRM). Instead of training a reward model, we request score feedback for each sentence from an advanced AI such as GPT-4 where represents previous sentences.
@@ 2.2 Reinforcement Learning
Reinforcement Learning (RL) fine-tunes our model to maximize sum of sentence rewards from an advanced AI model such as GPT-4 and Claude 3 Opus. The RL objective is as follows:
@@
Sentence-Level Nuanced Feedback: We request a score between and for each sentence in CoT through the advanced AI for RL. The closer the score is to , the more relevant and helpful it is to solving the problem. Table 5 presents the prompt format. We additionally shift the reward distribution by to center it at (Zheng et al., 2023). Therefore, the actual range is from to . Using these nuanced scores, the RL fine-tuned model exhibits emergent behaviors (please refer to Section 4). This allows us to understand the direction in which the model is intended to change through RL.
Advantages of Using Advanced AI for Score Feedback: Although calling the API has disadvantages, such as incurring costs or facing usage limits, there exist several advantages to using the advanced AI for feedback. First, there is no need to train a reward model. Second, as the RL fine-tuned model begins to generate out-of-distribution outputs that differ from the data used to train the reward model, it becomes challenging for the trained reward model to provide accurate rewards. However, this out-of-distribution problem is effectively addressed with the advanced AI.
@@ 2.3 Correction: Supervised Fine-Tuning
The RL fine-tuning procedure makes model changes to maximize the reward sum, such as correcting mistakes or explaining why other options cannot be the answer. However, without highly tuned hyperparameters (Eimer et al., 2023), the model after the RL phase may result in errors such as repeated sentences, truncated sentences, or incorrect content for some data points. (See examples in Appendix D.)
Correction Feedback: Given the success of LLMs and LMMs in a wide range of areas (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022a), we are not restricted to requesting feedback in the form of scores. We request correction feedback from advanced AI (Teacher) for sentences containing errors after the RL process, and obtain a corrected dataset . Since the supervised fine-tuning is more stable and finding appropriate hyperparameters is easier than RL, we proceed with supervised fine-tuning using exactly as in common autoregressive model (Vaswani et al., 2023) training to stabilize the RL fine-tuned model. This reduces the burden of RL’s exhaustive hyperparameter tuning and properly guides the direction in which the training model wants to change.
How Correction Feedback Helps RL: RL increases the probability of positively rewarded actions (or sentences) and decreases the probability for negative rewards. The direction of learning is determined by the reward (scalar) value. However, the opposite direction of the reward is sometimes required. For example, suppose there is a truncated sentence in CoT. gets a negative score because it is an incomplete sentence (Table 13). If there is no correction stage, the probability of is simply reduced. What if contains some valuable part? This valuable part is ignored, and its probability decreases. To alleviate this issue, we instead receive the corrected sentence as feedback and encourage the training model to generate complete sentences, which is very challenging to achieve with only RL. Table 16 shows more examples of how the correction stage helps the RL stage by maintaining the reasoning context while changing the erroneous parts.
@@
2.4 Algorithm Details
We propose a hybrid algorithm Alternating between REinforcement learning and Supervised fine-tuning (ARES). Figure 2 illustrates the ARES pipeline. First, we prepare a model with a given training dataset and generate rationale reasoning composed of several sentences for input. For the RL procedure to align the training model with a preference, we request scores for each sentence. The RL result may include some incorrect parts (colored in red as the 4th sentence in the RL result box), but it aims to maximize the rewards provided. Next, we request correction feedback and create a corrected dataset (colored in blue as the 3rd sentence in the Corrected Rationale box) for the supervised fine-tuning stage. We then repeat the process from the RL stage until convergence.
@@ 3 Experimental Setup
Baselines: We mainly compare our method with Multimodal-CoT (MM-CoT) (Zhang et al., 2023b) as the baseline because it utilizes reasoning chains to solve multi-modal tasks. MM-CoT leverages two distinct models: the first generates a rationale for a given problem, and the second, an inference model, takes the concatenated input (problem and generated rationale). This separated framework shows improved performance, even for relatively small models such as (Chia et al., 2023) (M) and (M). We use the rationale model provided by MM-CoT for ScienceQA and retrain the rationale model ourselves for A-OKVQA because there is no provided model.
Prompts for Feedback: Since our proposed ARES requests different types of feedback for each stage, a corresponding prompt exists separately. We use Claude 3 Haiku for all training to get feedback because it is approximately times cheaper than the top competing models, yet still demonstrates decent performance. We first request scores ranging from to for each sentence in CoT to proceed with the RL stage. To obtain reasonable scores, we let Haiku consider the starting point of thought, the process of elimination, or true statements. (See Table 5.)
@@
Training Details: For the RL stage, we use a learning rate of and epochs for PPO with a batch size of for both ScienceQA and A-OKVQA. The learning rate for is with epochs for PPO and a batch size of for both tasks. We proceed with 2 rounds of our pipeline for and 2 rounds for for ScienceQA. For A-OKVQA, we proceed with 1 round for both model sizes. For the SFT stage for correction, we follow the hyperparameters used in MM-CoT for both model sizes. Additionally, we replace MM-CoT’s inference model, which is the same size as the rationale model, with the Low-Rank Adaptation (LoRA) (Hu et al., 2021) added to the rationale model (Figure 4). The LoRA adapter effectively utilizes the rationale model’s features with a small number of weights, enabling 2x–14x faster inference compared to MM-CoT, which introduces a separate inference model (See the time comparison in Table 7 and Table 8). For more detailed settings, please refer to Appendix C.
@@ 4.2 Guide RL with Correction
Despite the benefits of RL, hyperparameter tuning often requires massive effort. Without meticulous tuning, the RL fine-tuned model may produce errors such as repetitive or incomplete sentences. To address these issues, we add a supervised fine-tuning (SFT) stage after RL to correct these errors. SFT is more stable than RL. We evaluate how well the SFT stage corrects errors caused by the RL stage for various RL hyperparameters. We test various RL hyperparameters such as learning rate = {5e-6, 1e-5, 2e-5, 5e-5}, batch size = {2, 4, 8, 16, 32}, and PPO epoch = {5, 10, 15}. As a result of RL, we observe that some of the sentences in rationale chains are repetitive or truncated (see Table 13 and 12). The SFT stage, with correction feedback, reflects the direction in which the model is fine-tuned through RL and appropriately guides it (Table 13 and 16). However, excessive RL learning rates or epochs cause serious degeneration of the model, such as producing no output or generating strange words, and the results of correction feedback are also unreasonable.
@@ 4.4 Inference Accuracy
We investigate whether the improved rationale also contributes to answer inference accuracy. Table 2 shows the main results of answer inference on the ScienceQA. We evaluate our base model against the MM-CoT baseline. achieves a 2.79% improvement compared to the corresponding baseline (). The large model () shows some minimal improvement compared to the corresponding baseline. However, it’s worth noting that despite this seemingly small gain, beats B LLaVA (Liu et al., 2023a). This minimal improvement may be due to the 9.5% of ScienceQA problems needing more rationale reasoning (around 9.5% problems have empty rationale reasoning). The RL stages can only eliminate some empty rationale reasoning, which requires numerous ARES pipeline rounds. Above all, our main goal is to assess how the RL stage works and how the SFT stage aids RL.
@@ 5 Related Work
Chain-of-Thought (CoT) is a multi-step reasoning method for problem-solving that encourages LLMs to consider the intermediate reasoning steps. Zero-Shot-CoT (Kojima et al., 2023) promotes CoT by using prompts such as "Let’s think step by step" for LLMs. For Few-Shot-CoT (Zhang et al., 2022b; Wei et al., 2023), a few examples with reasoning processes are provided, allowing the model to refer to these examples and understand how to perform CoT. Wei et al. (2023) reveal that this CoT technique positively impacts the performance of large models (B), but has minimal effect on smaller models. MM-CoT (Zhang et al., 2023b) suggest that CoT is beneficial even for relatively small models, such as 200M, if the model that generates intermediate reasoning and the model that infers the answer are separated. We find that simply adding a LoRA adapter (Hu et al., 2021) to the reasoning model results in comparable performance. This framework enables the LoRA adapter to effectively utilize all features, from raw text to latent features, and generates answers 2x–14x faster than MM-CoT, which uses a separate inference model (See Table 7 and Table 8). This speed advantage arises from the fact that our framework does not require a rationale as input, whereas the separate inference model framework must first generate the rationale before using it as input.